# Inductive-Deductive Strategy Reuse for Multi-Turn Instructional Dialogues

### v1

@@ Abstract
Aligning large language models (LLMs) with human expectations requires high-quality instructional dialogues, which can be achieved by raising diverse, in-depth, and insightful instructions that deepen interactions. Existing methods target instructions from real instruction dialogues as a learning goal and fine-tune a user simulator for posing instructions. However, the user simulator struggles to implicitly model complex dialogue flows and pose high-quality instructions. In this paper, we take inspiration from the cognitive abilities inherent in human learning and propose the explicit modeling of complex dialogue flows through instructional strategy reuse. Specifically, we first induce high-level strategies from various real instruction dialogues. These strategies are applied to new dialogue scenarios deductively, where the instructional strategies facilitate high-quality instructions. Experimental results show that our method can generate diverse, in-depth, and insightful instructions for a given dialogue history. The constructed multi-turn instructional dialogues can outperform competitive baselines on the downstream chat model.
@@ 1 Introduction
Large language models (LLMs) Du et al. (2022); OpenAI (2023) have demonstrated emergent capabilities across a wide range of language-related tasks naturally and interactively. This mainly benefits from the alignment of LLMs with humans. Aligning LLMs with humans requires high-quality instructional dialogues for fine-tuning LLMs. To automatically collect instructional dialogues, a feasible approach Xu et al. (2023b); Ding et al. (2023) is leveraging two top-tier LLMs to interact: one simulating a user to pose instructions, and the other acting as a system agent to respond. However, they primarily rely on LLMs conducting roleplay to simulate human behaviors via instructions, rather than learning from humans. This results in diminished diversity and insufficient depth of multi-turn instructional dialogues Kong et al. (2023b).
Some researchers have confirmed that raising high-quality instructions is a proactive way to make dialogues deeper and more diverse Yu et al. (2016); Li et al. (2017). Therefore, some works Kong et al. (2023b); Sun et al. (2023) target instructions from human-machine dialogues as a learning goal and fine-tune a user simulator to replace the role-played LLM. These methods assume that the user simulator can learn the shift trajectories of dialogues, i.e., dialogue flow Bao et al. (2023), without explicit guidance. However, the dialogue flow can take various directions for a given dialogue history Hou et al. (2018). The implicit modeling of complex dialogue flows leads to the user simulator tending to generate general instructions rather than diverse, in-depth, and insightful instructions.
@@
Figure 1: An example of humans generating instructions by deductively utilizing instructional strategies, derived from inductive analysis of instructional dialogues.
@@
In this paper, we propose to explicitly model dialogue flows to pose more diverse, in-depth, and insightful instructions. Inspired by the cognitive abilities in human learning Lin (1992), we observe that humans can skillfully utilize instructional strategies to promote the dialogue flow and generate high-quality instructions. Besides, humans can steer the dialogue flow towards different directions by choosing different strategies for generating diverse dialogues. Instructional strategies are general principles for guiding instruction generation, as shown in Figure 1. They are derived from inductive reasoning  Hayes et al. (2010) of various real dialogues, which can be applied to new dialogue scenarios via deductive reasoning Goel (2007).
@@
Motivated by this, we propose an Inductive-Deductive Strategy Reuse method, IDEAS for short, to generate diverse, in-depth, and insightful instructions for building multi-turn instructional dialogues. IDEAS is composed of an induction stage and a deduction stage. In the induction stage, we prompt GPT-4 to extract the instructional strategy for each pair of the dialogue history and the corresponding instruction in real human-machine dialogues. Those similar instructional strategies are further abstracted into a high-level strategy, which reveals the generic nature regardless of the details relevant to the specific dialogues. In the deduction stage, the user simulator is asked to select an appropriate instructional strategy to guide instruction generation based on the current dialogue history in each new dialogue scenario. Each selection is made by sampling one from a pool of potentially appropriate strategies to ensure diversity. Afterward, the user simulator and a system agent interact iteratively to build multi-turn instructional dialogues.
@@
Experiment results indicate that IDEAS can build high-quality instructional dialogues by generating diverse, in-depth, and insightful instructions. Moreover, our constructed instructional dialogues contribute to the performance improvement of the downstream chat model. Our contributions are summarized as follows: (1) We propose an inductive-deductive strategy reuse method to generate high-level instructions for new dialogue scenarios. To the best of our knowledge, this is the first study to pose more diverse, in-depth, and insightful instructions via strategy reuse. (2) We construct a high-quality multi-turn instructional dialogue dataset. (3) Experimental results show that IDEAS produces higher-quality instructions, which can be further used to improve the performance of the downstream chat model.
@@ 2 Preliminary
2.1 Task Definition
This work involves two tasks, including instruction generation and response generation. Please see their detailed definitions in Appendix A.
@@ 2.2 Induction & Deduction
Inductive reasoning, gaining common patterns and forming high-level principles from observations, is a core aspect of human intelligence Lake et al. (2017). Formally, inductive reasoning is to infer an unknown principle $\boldsymbol{f}:\boldsymbol{\mathcal{X}}\to\boldsymbol{\mathcal{Y}}$ that maps an input $\boldsymbol{x}\in\boldsymbol{\mathcal{X}}$ to an output $\boldsymbol{y}\in\boldsymbol{\mathcal{Y}}$. The principle $\boldsymbol{f}$ can take various forms, such as natural language descriptions Wang et al. (2023a). For instruction generation, inductive reasoning is inferring a principle $\boldsymbol{f}$ from several pairs of the dialogue histories $\{\boldsymbol{q}_{0},\boldsymbol{a}_{0},\dots,\boldsymbol{a}_{t-1}\}$ and the corresponding instructions $\boldsymbol{q}_{t}$. These dialogue histories often share common patterns and the principle $\boldsymbol{f}$ is prevalent in the mapping of all these pairs. In our study, we refer to the principle $\boldsymbol{f}$ as an instructional strategy.
@@
Deductive reasoning aims to derive new facts based on known facts and the induced principle Goel (2007). Formally, deductive reasoning is to derive an output $\boldsymbol{y}^{\prime}$ by applying $\boldsymbol{f}$ to a given input $\boldsymbol{x}^{\prime}$, i.e., $\boldsymbol{y}^{\prime}=\boldsymbol{f}(\boldsymbol{x}^{\prime})$. For instruction generation, deductive reasoning is deriving an instruction $\boldsymbol{q}^{\prime}_{t}$ based on a new dialogue history $\mathbf{h}^{\prime}_{t}=\{\boldsymbol{q}^{\prime}_{0},\boldsymbol{a}^{\prime}_{0},\dots,\boldsymbol{a}^{\prime}_{t-1}\}$ and an instructional strategy, i.e., the induced principle $\boldsymbol{f}$.
Figure 2: The overall architecture of building multi-turn instructional dialogues. In the induction stage, IDEAS induces high-level instructional strategies $\mathcal{F}$ from the real human-machine instructional dialogues $\mathcal{D}_{ins}$. In the deduction stage, the user simulator iteratively interacts with a system agent to produce new multi-turn dialogues based on the given opening line $\{\boldsymbol{q}^{\prime}_{0},\boldsymbol{a}^{\prime}_{0}\}$. The entire generation process is shown on the lower left side. For generating the current instruction $\boldsymbol{q}^{\prime}_{t}$, the user simulator first chooses an appropriate strategy $\boldsymbol{f}_{t}$ from the provided strategy candidate $Q(\boldsymbol{a}^{\prime}_{t-1})$ based on the dialogue history $\mathbf{h}^{\prime}_{t}$, and then generate an instruction $\boldsymbol{q}^{\prime}_{t}$. If the quality does not meet the requirement, $\boldsymbol{q}^{\prime}_{t}$ is regenerated. The corresponding process is shown on the lower right side.
@@ 3 IDEAS
In this section, we aim to reuse high-level instructional strategies to generate diverse, in-depth, and insightful instructions for building multi-turn instructional dialogues. To this end, we introduce inductive reasoning for gaining high-level strategies from human-machine instructional dialogues $\mathcal{D}_{ins}$ in Section 3.1. In section 3.2, we describe how the user simulator chooses an appropriate strategy to generate a high-quality instruction. In section 3.3, we show the training of models involved in Section 3.1 and 3.2. The overall architecture of building instructional dialogues is shown in Figure 2.
@@ 3.1 Induction Stage
This stage is to induce high-level instructional strategies $\mathcal{F}$ from the given human-machine dialogues $\mathcal{D}_{ins}=\{(\boldsymbol{q}_{0}^{i},\boldsymbol{a}_{0}^{i},\dots,\boldsymbol{q}_{T}^{i},\boldsymbol{a}_{T}^{i})_{i=1}^{N}\}$. Firstly, we successively extract each strategy expressed in natural language for each pair of dialogue history and the corresponding instruction. However, there are several challenges in reusing: (1) The instructional strategies may involve details related to specific dialogues, rendering their reuse challenging Zheng et al. (2024). (2) The myriad of strategies and their similarity further complicate the matter, as it is time-consuming to find the appropriate strategy for each dialogue history. Some examples are shown in Figure 2. Therefore, we conduct high-level abstractions on these instructional strategies to uncover their nature and reduce the strategy pool. Next, we delve into instructional strategy extraction and abstraction in detail.
Instructional Strategy Extraction.
For each pair of dialogue history $\mathbf{h}_{t}=\{\boldsymbol{q}_{0},\boldsymbol{a}_{0},\dots,\boldsymbol{a}_{t-1}\}$ and the corresponding instruction $\boldsymbol{q}_{t}$ from each dialogue in $\mathcal{D}_{ins}$, we provide GPT-4 with a task description and the input-output pair, and then prompt it to extract an instructional strategy $\boldsymbol{f}_{o}$. Specifically, we output $\boldsymbol{f}_{o}$ as follows:
@@
where $\boldsymbol{f}_{o}$ is expressed in natural language. This prompt can be found in Table 6. Consequently, we obtain original instructional strategies $\mathcal{F}_{o}$ on $\mathcal{D}_{ins}$.
Instructional Strategy Abstraction.
We observe that many instructional strategies bear similarities, they differ subtly in their expression or specific details. We thus generalize those similar strategies into a high-level strategy. A straightforward approach is to prompt GPT-4 to cluster original strategies $\mathcal{F}_{o}$ and then generalize for each cluster. However, it is unrealistic to feed all strategies into GPT-4 via a prompt due to the context length limitation Xu et al. (2023c). Besides, a single prompt cannot guide GPT-4 to handle complex tasks Li et al. (2023). Therefore, we perform high-level abstraction in two steps, including strategy clustering and cluster generalization.
For strategy clustering, we cannot approximate the number of clusters in advance. Thus, we inspire from Xu et al. (2023c) to form clusters by finding collections of similar instructional strategies. Each collection $\mathbf{C}_{i}$ contains the focused strategy $\boldsymbol{f}_{oi}$ and its similar strategies. Specifically, we retrieve similar strategies for each $\boldsymbol{f}_{oi}$ as follows:
@@
For cluster generalization, we also prompt GPT-4 to generate a high-level instructional strategy for each cluster. Please see Table 7 for this prompt. The size of each collection $\mathbf{C}_{i}$ is considerably smaller than $\mathcal{F}_{o}$, and it is relatively simple to generalize a set of similar strategies. Specifically, we provide GPT-4 with a generalization description and $\mathbf{C}_{i}$, and then output the high-level strategy $\boldsymbol{f}_{i}$ from the prompted GPT-4 as follows:
@@
Finally, the high-level instructional strategies =$\mathcal{F}=\{\boldsymbol{f}_{i}\}_{i=1}^{K}$ is applied to new dialogue scenarios.
@@ 3.2 Deduction Stage
This stage is to guide the user simulator to generate diverse, in-depth, and insightful instructions in a new dialogue scenario by reusing appropriate strategies in $\mathcal{F}$. We initiate a new dialogue scenario by providing a new opening line, which contains an instruction-answer pair $\{\boldsymbol{q}^{\prime}_{0},\boldsymbol{a}^{\prime}_{0}\}$. Afterward, the user simulator iteratively interacts with a system agent to produce new multi-turn dialogues under the given opening line. However, two sub-problems need to be addressed: (1) how to utilize the instructional strategy; and (2) how to ensure the quality of the generated instruction. Next, we respectively describe the corresponding solutions.
Instructional Strategy Utilization.
The natural idea is to leverage the chain-of-thought (CoT) technique Wei et al. (2022b), i.e., the user simulator first determines the appropriate instructional strategy based on the given dialogue history and then generate instructions. However, it is as challenging as generating a high-quality instruction directly Ou et al. (2022) since it also requires tracking and understanding the dialogue trajectories. Inspired by Zhu et al. (2023), we integrate a candidate set of strategies into the prompt, requiring the user simulator to retrieve a strategy from this set. If all strategies $\mathcal{F}$ are used as the candidate, the user simulator also suffers from the context length limitation. Thus, we roughly select $W$ potentially applicable strategies from $\mathcal{F}$ as the candidate.
Specifically, we introduce a ranker to calculate the probability that each $\boldsymbol{f}_{i}$ in $\mathcal{F}$ is suitable for the given dialogue history and consider those larger than a threshold $\eta$. We further randomly sample $W$ strategies as the candidate. More concretely, when generating the $t$-th instruction for the current dialogue history $\mathbf{h}^{\prime}_{t}=\{\boldsymbol{q}^{\prime}_{0},\boldsymbol{a}^{\prime}_{0},\dots,\boldsymbol{a}^{\prime}_{t-1}\}$ , the candidate is built as follows:
$Q(\boldsymbol{a}^{\prime}_{t-1})$ =$=\{\boldsymbol{f}_{i}|\operatorname{Ranker}(\boldsymbol{a}^{\prime}_{t-1},\boldsymbol{f}_{i})>\eta\},$ (4)
$Q(\boldsymbol{a}^{\prime}_{t-1})$ $=\operatorname{Sample}(Q(\boldsymbol{a}^{\prime}_{(t-1)}),W),$
@@
where $\boldsymbol{u}_{t}$ is a concatenated string with the form of [instructional strategy]$\boldsymbol{f}_{t}$[instruction]$\boldsymbol{q}^{\prime}_{t}$. This prompt can be found in Table 8. Overall, IDEAS can effectively model the dialogue flow by selecting an appropriate strategy, thereby generating an in-depth and insightful instruction. Sampling in Eq. (4) and (5) further ensures diversity.
@@
Specifically, we refer to widely-used metrics for dialogue evaluation Mehri and Eskenazi (2020), along with our goal, i.e., generating diverse, in-depth, and insightful instructions. We finally choose two appropriate metrics, correctness and coherence, to evaluate instructions. Correctness requires that the generated instruction $\boldsymbol{q}^{\prime}_{t}$ cannot contradict the dialogue history $\mathbf{h}^{\prime}_{t}$ and cannot be answered by existing answers $\{\boldsymbol{a}^{\prime}_{0},\dots,\boldsymbol{a}^{\prime}_{t-1}\}$, which can avoid talking about highly similar topics and promote the progress of the dialogue. Accordingly, coherence requires that $\boldsymbol{q}^{\prime}_{t}$ is related to $\mathbf{h}^{\prime}_{t}$, and coherently connected to the previous instructions or answers. The coherent instruction can lead to in-depth interaction. Since there are no automatic evaluation metrics available, we still prompt GPT-4 to judge instructions from these two dimensions. The corresponding prompt can be found in Table 9. Finally, we obtain the constructed multi-turn dialogues $\mathcal{D}_{chat}$ for training the chat model.
3.3 Model Training

### v2

@@ Abstract
Aligning large language models (LLMs) with human expectations requires high-quality instructional dialogues, which usually require instructions that are diverse and in-depth. Existing methods leverage two LLMs to interact for automatic collection: one simulating a user to pose instructions, and the other acting as a system agent to respond. However, these user simulators struggle to model the rules behind how dialogues can pose different instructions without explicit guidance, resulting in general instructions. In this paper, we propose to explicitly capture the complex rules to help the user simulator pose diverse and in-depth instruction. Specifically, we first induce high-level instruction strategies from various real instruction dialogues serving as rules. Afterward, different possible strategies are applied to the newly given dialogue scenario deductively to pose various instructions. Experimental results show that our method can generate diverse and in-depth instructions. The constructed multi-turn instructional dialogues can outperform competitive baselines on the downstream chat model  [^1] https://github.com/kwai/IDEAS.
@@ 1 Introduction
Large language models (LLMs) Du et al. (2022); OpenAI (2023) have demonstrated emergent capabilities across a wide range of language-related tasks naturally and interactively. This mainly benefits from the alignment of LLMs with humans. The alignment requires high-quality multi-turn instructions for fine-tuning LLMs in a dialogue-based setting Wang et al. (2023c). Such instructional dialogues usually require instructions that are diverse and in-depth Yu et al. (2016); Li et al. (2017). Specifically, diverse instructions help instruction-following LLMs adapt to different situations and user needs. In-depth instructions help LLMs follow the logical flows of dialogues and deepen the understanding of dialogue histories to answer users’ instructions. However, manually collecting such instructional dialogue data is usually labor-intensive and time-consuming Wang et al. (2023b).
@@
Figure 1: An example of humans generating instructions by deductively utilizing instruction strategies, derived from inductive analysis of instructional dialogues.
A feasible approach to automatic collection is leveraging two LLMs to interact: one simulating a user to pose instructions, and the other acting as a system agent to respond. User simulators are currently implemented by employing role-played LLMs to simulate humans through instruction prompts Xu et al. (2023b); Ding et al. (2023) or fine-tuning LLMs to learn instruction generation from real instructional dialogues Kong et al. (2023b); Sun et al. (2023). However, dialogue data allows for different dialogue flows for the given dialogue history, i.e., various instructions can be posed Hou et al. (2018). LLMs in these user simulators implicitly learn the underlying rules guiding these different flows from training data. Due to the complexity and variability of these rules, LLMs struggle to capture various possible rules without explicit guidance so that the user simulators tend to capture frequently co-occurring patterns and generate more generic instructions Li et al. (2017). This results in diminished diversity and insufficient depth of multi-turn instructions.
@@
In this paper, we propose to explicitly model the rules behind how dialogues can flow in different directions, posing diverse and in-depth instructions. Inspired by the cognitive abilities in human learning Lin (1992), we observe that humans can gain general rules from various real dialogues via inductive reasoning  Hayes et al. (2010), which guides the dialogue flows in certain directions. As shown in Figure 1, humans induce rules such as “Ask for specific information to clarify detail” and “Information validation and illustration”. We refer to these rules as instruction strategies. Afterward, humans can pose various instructions in new dialogue scenarios by choosing different possible strategies via deductive reasoning Goel (2007). Specifically, different strategies focus on different aspects, such as the “detail” or “correctness” of information, which directs the dialogue flows in various directions.
@@
Motivated by this, we propose an Inductive-Deductive Strategy Reuse method, IDEAS for short, to generate diverse and in-depth instructions for building multi-turn instructional dialogues. IDEAS is composed of an induction stage and a deduction stage. In the induction stage, we prompt GPT-4 to extract the instruction strategy for each pair of the dialogue history and the corresponding instruction in real human-machine dialogues. Those similar instruction strategies are further abstracted into a high-level strategy, which reveals the generic nature regardless of the details relevant to the specific dialogues. In the deduction stage, the user simulator is asked to sample an appropriate instruction strategy to guide instruction generation based on the current dialogue history in each new dialogue scenario. Afterward, the user simulator and a system agent interact iteratively to build multi-turn instructional dialogues.
@@
Experiment results indicate that IDEAS can generate diverse and in-depth instructions, and our constructed multi-turn instructional dialogues contribute to the performance improvement of the downstream chat model. Our contributions are summarized as follows: (1) To the best of our knowledge, this is the first study to propose an inductive-deductive strategy reuse method to generate diverse and in-depth instructions in new dialogue scenarios. (2) Experimental results show that IDEAS produces higher-quality instructions, which can be further used to improve the performance of the downstream chat models. (3) Extensive experiments show that providing more instructional dialogues with high-quality instructions can further improve performance.
@@ 2 Preliminary
2.1 Task Definitions
In this paper, we first study the generation of high-quality instructions to construct multi-turn instructional dialogues and then use the constructed instruction dialogues to fine-tune the chat model to improve the model’s ability to answer instructions. Thus, this work involves two tasks, including instruction generation and answer generation.
Instruction Generation.
This generative model can be a role-played LLM or a fine-tuned LLM on real human-machine dialogues as a user simulator, which iteratively interacts with a system agent (e.g., GPT-4) to obtain multi-turn instructional dialogues for different dialogue scenarios.
Answer Generation.
@@ 2.2 Induction & Deduction
Inductive reasoning, gaining common patterns and forming high-level rules from observations, is a core aspect of human intelligence Lake et al. (2017). Formally, inductive reasoning is to infer an unknown rule $\boldsymbol{f}:\boldsymbol{\mathcal{X}}\to\boldsymbol{\mathcal{Y}}$ that maps an input $\boldsymbol{x}\in\boldsymbol{\mathcal{X}}$ to an output $\boldsymbol{y}\in\boldsymbol{\mathcal{Y}}$. The rule $\boldsymbol{f}$ can take various forms, such as natural language descriptions Wang et al. (2023a). For instruction generation, inductive reasoning is inferring an instruction strategy $\boldsymbol{f}$ from several pairs of the dialogue histories $\{\boldsymbol{q}_{0},\boldsymbol{a}_{0},\dots,\boldsymbol{a}_{t-1}\}$ and the corresponding instructions $\boldsymbol{q}_{t}$. These dialogue histories often share common patterns and the rule $\boldsymbol{f}$ is prevalent in the mapping of all these pairs.
@@
Deductive reasoning aims to derive new facts based on known facts and the induced rules Goel (2007). Formally, deductive reasoning is to derive an output $\boldsymbol{y}^{\prime}$ by applying $\boldsymbol{f}$ to a given input $\boldsymbol{x}^{\prime}$, i.e., $\boldsymbol{y}^{\prime}=\boldsymbol{f}(\boldsymbol{x}^{\prime})$. For instruction generation, deductive reasoning is deriving an instruction $\boldsymbol{q}^{\prime}_{t}$ based on a new dialogue history $\mathbf{h}^{\prime}_{t}=\{\boldsymbol{q}^{\prime}_{0},\boldsymbol{a}^{\prime}_{0},\dots,\boldsymbol{a}^{\prime}_{t-1}\}$ and an instruction strategy, i.e., the induced rule $\boldsymbol{f}$.
Figure 2: The overall architecture of building multi-turn instructional dialogues. In the induction stage, IDEAS induces high-level strategies $\mathcal{F}$ from human-machine instructional dialogues $\mathcal{D}_{ins}$. In the deduction stage, the user simulator iteratively interacts with a system agent to produce new multi-turn dialogues based on the given opening line $\{\boldsymbol{q}^{\prime}_{0},\boldsymbol{a}^{\prime}_{0}\}$, as shown on the left side. For generating the current instruction $\boldsymbol{q}^{\prime}_{t}$ , the user simulator first chooses an appropriate strategy $\boldsymbol{f}_{t}$ from the candidate $Q(\boldsymbol{a}^{\prime}_{t-1})$ based on the dialogue history $\mathbf{h}^{\prime}_{t}$, and then generate $\boldsymbol{q}^{\prime}_{t}$. If the quality does not meet the requirement, $\boldsymbol{q}^{\prime}_{t}$ is regenerated. This process is shown on the right side.
@@ 3 IDEAS
We aim to reuse instruction strategies to generate diverse and in-depth instructions for building multi-turn instructional dialogues. To this end, we introduce inductive reasoning for gaining high-level strategies from human-machine instructional dialogues $\mathcal{D}_{ins}$ in Section 3.1. In Section 3.2, we describe how the user simulator chooses an appropriate strategy to generate instruction. In Section 3.3, we show the implementation of the involved models. The overall architecture of building instructional dialogues is shown in Figure 2.
@@ 3.1 Induction Stage
This stage is to induce instruction strategies $\mathcal{F}$ from the given human-machine dialogues $\mathcal{D}_{ins}=\{(\boldsymbol{q}_{0}^{i},\boldsymbol{a}_{0}^{i},\dots,\boldsymbol{q}_{T}^{i},\boldsymbol{a}_{T}^{i})_{i=1}^{N}\}$. Firstly, we successively extract each strategy expressed in natural language for each pair of dialogue history and the corresponding instruction. However, there are several challenges in reusing: (1) The strategies may involve details related to specific dialogues, rendering their reuse challenging Zheng et al. (2024a). (2) The variability and similarity of strategies further complicate the matter, as it is time-consuming to find the appropriate strategy for each dialogue history. Some examples are shown in Figure 2. Therefore, we conduct high-level abstractions on these strategies to uncover their nature and reduce the strategy pool. Next, we delve into instruction strategy extraction and abstraction in detail.
Instruction Strategy Extraction.
For each pair of dialogue history $\mathbf{h}_{t}=\{\boldsymbol{q}_{0},\boldsymbol{a}_{0},\dots,\boldsymbol{a}_{t-1}\}$ and the corresponding instruction $\boldsymbol{q}_{t}$ from each dialogue in $\mathcal{D}_{ins}$, we prompt GPT-4 with a task description and the input-output pair to extract an instruction strategy $\boldsymbol{f}_{o}$ as follows:
@@
where $\boldsymbol{f}_{o}$ is expressed in natural language. This prompt can be found in Table 9. Consequently, we obtain original instruction strategies $\mathcal{F}_{o}$ on $\mathcal{D}_{ins}$.
Instruction Strategy Abstraction.
We observe that many strategies bear similarities, they differ subtly in their expression or specific details. We thus generalize those similar strategies into a high-level strategy. A straightforward approach is to prompt GPT-4 to cluster original strategies $\mathcal{F}_{o}$ and then generalize for each cluster. However, it is unrealistic to feed all strategies into GPT-4 due to the context length limitation Xu et al. (2023c). Besides, a single prompt cannot guide GPT-4 to handle complex tasks Li et al. (2023). Thus, we perform high-level abstraction in two steps, including strategy clustering and cluster generalization.
For strategy clustering, we cannot approximate the number of clusters in advance. Thus, we take inspiration from Xu et al. (2023c) to form clusters by finding collections of similar strategies. Each collection $\mathbf{C}_{i}$ contains the focused strategy $\boldsymbol{f}_{oi}$ and its similar strategies. Specifically, we take each original strategy $\boldsymbol{f}_{oi}$ as the focused strategy in turn and retrieve similar strategies for each $\boldsymbol{f}_{oi}$ as follows:
@@
For cluster generalization, we also prompt GPT-4 to generate a high-level instruction strategy for each cluster. Please see Table 10 for this prompt. The size of each collection $\mathbf{C}_{i}$ is considerably smaller than $\mathcal{F}_{o}$, and it is relatively simple to generalize a set of similar strategies. Specifically, we prompt GPT-4 with task description and $\mathbf{C}_{i}$, and then output the high-level strategy $\boldsymbol{f}_{i}$ as follows:
@@
Finally, the high-level instruction strategies =$\mathcal{F}=\{\boldsymbol{f}_{i}\}_{i=1}^{K}$ are applied to new dialogue scenarios.
@@ 3.2 Deduction Stage
This stage is to guide the user simulator to generate diverse and in-depth instructions in a new dialogue scenario by reusing appropriate strategies in $\mathcal{F}$. We initiate a new dialogue scenario by providing a new opening line, which contains an instruction-answer pair $\{\boldsymbol{q}^{\prime}_{0},\boldsymbol{a}^{\prime}_{0}\}$. Afterward, the user simulator iteratively interacts with a system agent to produce new multi-turn dialogues under the given opening line. However, two sub-problems need to be addressed: (1) how to utilize the strategy; and (2) how to ensure the quality of the generated instruction. Next, we describe the corresponding solutions.
Instruction Strategy Utilization.
The natural idea is to leverage the chain-of-thought (CoT) technique Wei et al. (2022b), i.e., the user simulator first determines the appropriate instruction strategy based on the given dialogue history and then generates instructions. However, it is as challenging as generating high-quality instructions directly since it also requires tracking and understanding the dialogue flows Ou et al. (2022). Inspired by Zhu et al. (2023), we provide a candidate set of instruction strategies, requiring the user simulator to retrieve a strategy from this set. If all strategies $\mathcal{F}$ are used as the candidate, the user simulator also suffers from the context length limitation. Thus, we roughly select $W$ potentially applicable strategies from $\mathcal{F}$ as the candidate.
Specifically, we introduce a ranker to calculate the probability that each instruction strategy $\boldsymbol{f}_{i}$ in $\mathcal{F}$ is suitable for the given dialogue history and consider those larger than a threshold $\eta$. We further randomly sample $W$ strategies as the candidate. More concretely, when generating the $t$-th instruction for the current dialogue history $\mathbf{h}^{\prime}_{t}=\{\boldsymbol{q}^{\prime}_{0},\boldsymbol{a}^{\prime}_{0},\dots,\boldsymbol{a}^{\prime}_{t-1}\}$ , the candidate $Q(\boldsymbol{a}^{\prime}_{t-1})$ is built as follows:
$Q^{\prime}(\boldsymbol{a}^{\prime}_{t-1})$$\leftarrow\{\boldsymbol{f}_{i}|\operatorname{Ranker}(\boldsymbol{a}^{\prime}_{t-1},\boldsymbol{f}_{i})>\eta\},$ (6)
$Q(\boldsymbol{a}^{\prime}_{t-1})$ $\leftarrow\operatorname{Sample}(Q^{\prime}(\boldsymbol{a}^{\prime}_{(t-1)}),W),$
@@
where $\boldsymbol{u}_{t}$ is a concatenated string with the form of [instruction strategy]$\boldsymbol{f}_{t}$[instruction]$\boldsymbol{q}^{\prime}_{t}$. This prompt can be found in Table 11.
@@
Specifically, we refer to widely-used metrics for dialogue evaluation Mehri and Eskenazi (2020), along with our goal, i.e., generating diverse and in-depth instructions. We finally choose two appropriate metrics, correctness and coherence, to evaluate instructions. Correctness requires that the generated instruction $\boldsymbol{q}^{\prime}_{t}$ cannot contradict the dialogue history $\mathbf{h}^{\prime}_{t}$ and cannot be answered by existing answers $\{\boldsymbol{a}^{\prime}_{0},\dots,\boldsymbol{a}^{\prime}_{t-1}\}$, which can avoid talking about highly similar topics and promote the progress of the dialogue. Accordingly, coherence requires that $\boldsymbol{q}^{\prime}_{t}$ is related to $\mathbf{h}^{\prime}_{t}$, and coherently connected to the previous instructions or answers. The coherent instruction can lead to in-depth interaction. Since there are no automatic evaluation metrics available, we still prompt GPT-4 to judge instructions from these two dimensions. The corresponding prompt can be found in Table 12. Finally, we obtain the constructed multi-turn dialogues $\mathcal{D}_{chat}$ for training the chat model. Please refer to B.2 for the corresponding training details.
3.3 Model Implementation